# On complexity of restricted fragments of Decision DNNF

### v1

@@ 2.2 Models of Boolean functions
We assume the reader to be familiar with the notion of Boolean formulas and in particular, with the notion of conjunctive normal forms (cnfs). For a cnf $\varphi$, $\mathsf{var}(\varphi)$ is the set of variables whose literals occur in the clauses of $\varphi$. For example, let $\varphi=(x_{1}\vee x_{2})\wedge(x_{2}\vee\neg x_{3}\vee\neg x_{4})$. $\mathsf{var}(\varphi)=\{x_{1},x_{2},x_{3},x_{4}\}$.
In order to define the function represented by a cnf, it will be convenient to consider a cnf as a set of its clauses and a clause as a set of literals. Let ${\bf a}$ be an assignment and let $L_{\bf a}$ be the well-formed set of literals over $\mathsf{var}({\bf a})$ corresponding to ${\bf a}$. Let $\varphi=\{C_{1},\dots,C_{q}\}$ be a cnf. Then ${\bf a}$ satisfies a clause $C_{i}$ if $L_{\bf a}\cap C_{i}\neq\emptyset$ and ${\bf a}$ satisfies $\varphi$ if ${\bf a}$ satisfies each clause of $\varphi$. The function $f(\varphi)$ represented by $\varphi$ has the set of variables $\mathsf{var}(\varphi)$ and $\mathcal{S}(f(\varphi))$ consists of exactly those assignments that satisfy $\varphi$. In what follows, we slightly abuse the notation and refer to the function represented by $\varphi$ also by $\varphi$. The correct use will always be clear from the context.
We are next going to define the model Decision dnnf, which throughout this paper is referred to as $\wedge_{d}$-fbdd (the reasons will become clear later). The model is based on a directed acyclic graph (dag) with a single source with vertices and some edges possibly being labelled. We will often use the following notation. Let $D$ be a dag and let $u$ be a vertex of $D$. Then we denote by $D_{u}$ the subgraph of $D$ induced by $u$ and all the vertices reachable from $u$ in $D$. The labels on vertices and edges of $D_{u}$ are the same as the labels of the respective vertices and edges of $D$.
@@
A $\wedge_{d}$-fbdd is a dag $B$ with a single source and two sinks. One sink is labelled by ${\bf 0}$, the other by ${\bf 1}$. Each non-sink node has two out-neighbours (children). A non-sink node may be either a conjunction node, labelled with $\wedge$ or a decision node labelled with a variable name. If $u$ is a decision node then one outgoing edge is labelled with 00, the other with $1$.
Let $u$ be a node of $B$. Then $\mathsf{var}(B_{u})$ is the set of variables labelling the decision nodes of $B_{u}$. $B$ obeys two constraints: decomposability and read-onceness. The decomposability constraint means that all the conjunction nodes $u$ are decomposable. In particular, let $u_{0}$ and $u_{1}$ be the children of $u$. Then $\mathsf{var}(B_{u_{0}})\cap\mathsf{var}(B_{u_{1}})=\emptyset$. The read-onceness means that two decision nodes labelled by the same variable cannot be connected by a directed path.
@@
Let $B$ be a $\wedge_{d}$-fbdd. We define the set $\mathcal{A}(B)$ of assignments accepted by $B$ in the following recursive way. If $B$ consists of a sink node only then $\mathcal{A}(B)=\{\emptyset\}$ if the sink is labelled by ${\bf 1}$ and $\emptyset$ otherwise.
@@
Assume first that $u$ is a decision node and let $x$ be the variable labelling $u$. We assume w.l.o.g that $(u,u_{0})$ is labelled by 00 and $(u,u_{1})$ is labelled by $1$. Then $\mathcal{A}(B)=\mathcal{A}_{0}\times\{(x,0)\}\cup\mathcal{A}_{1}\times\{(x,1)\}$. If $u$ is a conjunction node then $\mathcal{A}(B)=\mathcal{A}_{0}\times\mathcal{A}_{1}$.
The function $f(B)$ represented by $B$ is the function over $var(B)$ whose satisfying assignments are all ${\bf b}$ such that there is ${\bf a}\in\mathcal{A}(B)$ such that ${\bf a}\subseteq{\bf b}$. We denote $\mathcal{S}(f(B))$ by $\mathcal{S}(B)$ and refer to this set as the set of satisfying assignments of $B$.
@@
We next define subclasses of $\wedge_{d}$-fbdds. Let $B$ be a $\wedge_{d}$-fbdd that does not have $\wedge_{d}$ nodes. Then $B$ is a fbdd. Next, let $B$ be a $\wedge_{d}$-fbdd such that there is a linear order $\pi$ over a superset of $\mathsf{var}(B)$ so that the following holds. Let $P$ be a directed path of $B$ and suppose that $P$ has two decision nodes $u_{1}$ and $u_{2}$ labelled by variables $x_{1}$ and $x_{2}$ and assume that $u_{1}$ occurs before $u_{2}$ in $P$. Then $x_{1}<_{\pi}x_{2}$ (that is $x_{1}$ is smaller than $x_{2}$ according to $\pi$). Then $B$ is called an $\wedge_{d}$-obdd (obeying $\pi$ if not clear from the context). Finally, an $\wedge_{d}$-obdd without $\wedge_{d}$ nodes is just an obdd. Allowing $\pi$ to be defined over a superset of $\mathsf{var}(B)$ will be convenient for our reasoning when we consider subgraphs of a $\wedge_{d}$-obdd.
@@
Throughout the paper, when we refer to a smallest model representing a function $f$, we will use the notation $modelname(f)$. For instance $\wedge_{d}$-obdd$(f)$, $(f)$, obdd$(f)$ respectively refer to the smallest $\wedge_{d}$-obdd, fbdd($f$), and obdd$(f)$ representing $f$.
@@ 2.3 Graphs, width parameters, and correspondence with cnf
We use the standard terminology related to graph as specified in [7]. Let us recall that a matching $M$ of a graph $G$ is a subset of its edges such that for any distinct $e_{1},e_{2}\in M$, $e_{1}\cap e_{2}=\emptyset$. The treewidth of $G$, denoted by $tw(G)$ is arguably the best known graph parameter measuring closeness to a tree. To define $tw(G)$, we need the notion of a tree decomposition of $G$ which is a pair $(T,{\bf B})$ where $T$ is a tree and ${\bf B}:V(T)\rightarrow 2^{V(G)}$. The set ${\bf B}(t)$ for $t\in V(T)$ is called the bag of $T$. The bags obey the rules of containement meaning that for each $e\in V(G)$ there is $t\in V(T)$ such that $e\subseteq{\bf B}(t)$ and connectivity. For the latter, let $T[v]$ be the subgraph of $T$ induced by $t$ such that $v\in{\bf B}(t)$. Then $T[v]$ is connected. For $t\in V(T)$, the width of ${\bf B}(t)$ is $|{\bf B}(t)|-1$ The width of $(T,{\bf B})$ is the largest width of a bag. Finally, $tw(G)$ is the smallest width of a tree decomposition. The pathwidth of $G$, denoted by $pw(G)$, is the smallest width of a tree decomposition $(T,{\bf B})$ of $G$ such that $T$ is a path.
@@
For a cnf $\varphi$, we use two graphs, both regard $\varphi$ as a hypergraph over $\mathsf{var}(\varphi)$ ignoring the literals. The primal graph of $\varphi$ has $\mathsf{var}(\varphi)$ as the set of vertices and the two vertices are adjacent if they occur in the same clause. The incidence graph of $\varphi$ has $\mathsf{var}(\varphi)$ and the clauses as its vertices. A variable $x$ is adjacent to a clause $C$ if and only if $x$ occurs in $C$. We denote by $ptw(\varphi)$ and $itv(\varphi)$ the trewidth of the primal and incidence graphs of $\varphi$ respectively.
@@
3 $\wedge_{d}$-obdd: incidence treewidth inefficiency and separations
3.1 An approach to proving $\wedge_{d}$-obbd lower bounds
We start from overviewing an approach for proving lower bounds for obdds and then generalize it to $\wedge_{d}$-obdds.
Let $B$ be an obdd obeying an order $\pi$. Let ${\bf g}$ be an assigment over a prefix of $\pi$. Then we can identify the maximal path $P=P({\bf g})$ starting from the source and such that ${\bf a}(P)\subseteq{\bf g}$. Indeed, existence of such a path and its uniqueness is easy to verify algorithmically. Let the current vertex $u$ of $B$ be the source and the current variable $x$ be the label of $u$. Furthermore, let the current path $P$ be consisting of $u$ only. If $x\notin\mathsf{var}({\bf g})$ then stop. Otherwise, choose the out-neighbour $v$ of $u$ such that $(u,v)$ is labelled with ${\bf g}(x)$. Let the current vertex be $v$, the current variable be the label of $v$ and the current path be $P$ plus $(u,v)$. Repeat the process until the stopping condition reached. We let $u({\bf g})$ be the final vertex of $P({\bf g})$.
Now, let $\mathcal{F}$ be a set of fooling assignments, for each ${\bf g}\in\mathcal{F}$, $Var({\bf g})$ is a prefix of $P$. If we prove that for every two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$ then this implies that $|B|\geq|\mathcal{F}|$. In particular, if $|\mathcal{F}|$ is large in terms of the number of variables, then the respective large lower bound follows.
If we try to adapt the above methodology to a $\wedge_{d}$-obdd, $B$ obeying an order $\pi$, we will notice that an assignment ${\bf g}$ as above is associated not with a path but rather with a rooted tree. The role is $u({\bf g})$ is played by the leaves of that tree. Consequently, there is no straightforward way to identify a single node associated with an assignment. We propose a generic methodology that helps to identify such a single node. In the rest of this subsection, we provide a formal description of this methodology.
@@ 3.1 An approach to proving $\wedge_{d}$-obbd lower bounds
Figure 1 demonstrates a $\wedge_{d}$-obdd $B$ for $(x_{1}\vee x_{2})\wedge(x_{2}\vee x_{3})\wedge(x_{4}\vee x_{5})\wedge(x_{5}\vee x_{6})$ (on the left) obeying the order $(x_{2},x_{1},x_{3},x_{5},x_{4},x_{6})$ and $B[{\bf g}]$ (on the right) where ${\bf g}=\{(x_{2},1)\}$.
@@
A decision node of $B[{\bf g}]$ that has only one outgoing neighbour is called an incomplete decision node. A maximal path all intermediate decision nodes of which are incomplete is called an incomplete path. We note that a $\wedge_{d}$ node cannot be a final node of an incoplete path due to its maximality, hence an incomplete path may end only with a sink of $B$ or with a decision node of $B$. We denote by $L({\bf g})$ the set of all decision nodes that are final nodes of the incomplete paths of $B[{\bf g}]$.
Back to Example 2, $L({\bf g})$ is a singleton consisting of the only node labelled by $x_{5}$. We note that $L({\bf g})$ are exactly minimal complete decision nodes of $B[{\bf g}]$.
@@
Let $\pi$ be the order obeyed by $B$ containing $\mathsf{var}(B)$ and let ${\bf g}$ be an assignment over a prefix of $\pi$. Then the following statements hold regarding $B[{\bf g}]$.
@@
    Let $T({\bf g})$ be the union of all incomplete paths ending with nodes of $L({\bf g})$. Then $T({\bf g})$ is a rooted tree.
@@
Let $x$ and $y$ be the variables labelling $u$ and $v$ respectively. By assumption $x<_{\pi}y$ and $x\notin Var({\bf g})$, the latter constitutes a prefix of $\pi$. It follows that $y\notin Var({\bf g})$ and hence $v$ is complete simply by construction.
@@
For the third statement, it is sufficient to prove that any two distinct incomplete paths $P_{1}$ and $P_{2}$ do not have nodes in common besides their comon prefixes. So, let $P$ be the longest common prefix of $P_{1}$ and $P_{2}$ and let $u$ be the last vertex of $P_{1}$ and $P_{2}$. Arguing as in the previous paragraph, we conclude that $u$ is a $\wedge_{d}$ node. Let $u_{1}$ and $u_{2}$ be the children of $u$. Then, by the maximality of $P$, we can assume w.l.o.g. that $u_{1}\in V(P_{1})$ and $u_{2}\in V(P_{2})$.
@@
Following Lemma 2, we denote by $T({\bf g})$ the union of incomplete paths of $B[{\bf g}]$ ending with a node of $L({\bf g})$. The tree $T({\bf g})$ will not be essential for our further reasoning. However, it is important to underscore the direction of upgrade reasoning as we move from obddto $\wedge_{d}$-obdd: instead of the path $P({\bf g})$, we consider the tree $T({\bf g})$.
With the premises of Lemma 2, $L({\bf g})$ provides a neat characterization of $\mathcal{S}(B)|_{\bf g}$ as established in the next lemma.

### v2

@@ 2.2 Models of Boolean functions
We assume the reader to be familiar with the notion of Boolean formulas and in particular, with the notion of conjunctive normal forms (cnfs). For a cnf $\varphi$, $\mathsf{var}(\varphi)$ is the set of variables whose literals occur in the clauses of $\varphi$. For example, let $\varphi=(x_{1}\vee x_{2})\wedge(x_{2}\vee\neg x_{3}\vee\neg x_{4})$. Then $\mathsf{var}(\varphi)=\{x_{1},x_{2},x_{3},x_{4}\}$.
In order to define the function represented by a cnf, it will be convenient to consider a cnf as a set of its clauses and a clause as a set of literals. Let $\bf a$} be an assignment and let $L_{\bf a}$ be the well-formed set of literals over $\mathsf{var}({\bf a})$ corresponding to $\bf a$}. Let $\varphi=\{C_{1},\dots,C_{q}\}$ be a cnf. Then $\bf a$} satisfies a clause $C_{i}$ if $L_{\bf a}\cap C_{i}\neq\emptyset$ and $\bf a$} satisfies $\varphi$ if $\bf a$} satisfies each clause of $\varphi$. The function $f(\varphi)$ represented by $\varphi$ has the set of variables $\mathsf{var}(\varphi)$ and $\mathcal{S}(f(\varphi))$ consists of exactly those assignments (over $\mathsf{var}(\varphi)$) that satisfy $\varphi$. In what follows, we slightly abuse the notation and refer to the function represented by $\varphi$ also by $\varphi$. The correct use will always be clear from the context.
Remark 1.
Let us remark that, for a cnf $\varphi$, $\mathcal{S}(\varphi)$ consists of the satisfying assignment of $\varphi$ over all the variables of $\varphi$ (simplyt because, in this context we consider $\varphi$ a function). On the other if $\varphi$ is viewed as a cnf, it can have much shorter satisfying assignments. For example, $\varphi=(x_{1}\vee x_{2})\wedge(x_{2}\vee\neg x_{3}\vee\neg x_{4})$ is satisfied with $\{(x_{2},1)\}$. It is well known that $\mathcal{S}(\varphi)$ is the set of all extensions (to the whole of $\mathsf{var}(\varphi)$) of all the minimal satisfying assignments of $\varphi$.
We are next going to define the model Decision dnnf, which throughout this paper is referred to as $\wedge_{d}$-fbdd (the reasons will become clear later). The model is based on a directed acyclic graph (dag) with a single source with vertices and some edges possibly being labelled. We will often use the following notation. Let $D$ be a dag. Then $V(D)$ is the set of vertices (nodes) of $D$. Further on, let $u\in V(D)$. Then we denote by $D_{u}$ the subgraph of $D$ induced by $u$ and all the vertices reachable from $u$ in $D$. The labels on vertices and edges of $D_{u}$ are the same as the labels of the respective vertices and edges of $D$.
@@
A $\wedge_{d}$-fbdd is a dag $B$ with a single source and two sinks. One sink is labelled by $True$, the other by $False$ (on the illustrations we denote the sinks by $T$ and $F$ respectively). Each non-sink node has two out-neighbours (children). A non-sink node may be either a decomposable conjunction node, labelled with $\wedge_{d}$ (we often refer to such nodes as $\wedge_{d}$-nodes) or a decision node labelled with a variable name. If $u$ is a decision node then one outgoing edge is labelled with 0, the other with $1$.
Let $u$ be a node of $B$. Then $\mathsf{var}(B_{u})$ is the set of variables labelling the decision nodes of $_{u}. BB$ obeys two constraints: decomposability and read-onceness. The decomposability constraint means that if $u$ is a $\wedge_{d}$-node and $u_{0}$ and $u_{1}$ are the children of $u$ then $\mathsf{var}(B_{u_{0}})\cap\mathsf{var}(B_{u_{1}})=\emptyset$. The read-onceness means that two decision nodes labelled by the same variable cannot be connected by a directed path.
@@
Let $B$ be a $\wedge_{d}$-fbdd. We define the set $\mathcal{A}(B)$ of assignments accepted by $B$ in the following recursive way. If $B$ consists of a sink node only then $\mathcal{A}(B)=\{\emptyset\}$ if the sink is labelled by $True$ and $\emptyset$ otherwise.
@@
Assume first that $u$ is a decision node and let $x$ be the variable labelling $u$. We assume w.l.o.g that $(u,u_{0})$ is labelled by 0 and $(u,u_{1})$ is labelled by $1$. Then $\mathcal{A}(B)=\mathcal{A}_{0}\times\{(x,0)\}\cup\mathcal{A}_{1}\times\{(x,1)\}$. If $u$ is a conjunction node then $\mathcal{A}(B)=\mathcal{A}_{0}\times\mathcal{A}_{1}$.
The function $f(B)$ represented by $B$ is the function over $\mathsf{var}(B)$ whose satisfying assignments are all $\bf b$} such that there is ${\bf a}\in\mathcal{A}(B)$ such that ${\bf a}\subseteq{\bf b}$. We denote $\mathcal{S}(f(B))$ by $\mathcal{S}(B)$ and refer to this set as the set of satisfying assignments of $B$.
Example 1.
Figure 1 illustrates $\wedge_{d}$-fbdd $B$ representing function $(x\rightarrow(y_{1}\wedge y_{2}))\vee(\neg x\rightarrow(y_{1}\wedge y_{3}))$. In particular, $\mathcal{A}(B)=\{\{(x,1),(y_{1},1),(y_{2},1)\},\{(x,0),(y_{1},1),(y_{3},1)\}\}$. Clearly, $\mathcal{S}(B)$ are all possible supersets of the assignments of $\mathcal{A}(B)$ assigning all of $x,y_{1},y_{2},y_{3}$.
[…]
Figure 1: An example of a $\wedge_{d}$-obdd
@@
We next define subclasses of $\wedge_{d}$-fbdds. Let $B$ be a $\wedge_{d}$-fbdd that does not have $\wedge_{d}$ nodes. Then $B$ is a fbdd. Next, let $B$ be a $\wedge_{d}$-fbdd such that there is a linear order $\pi$ over a superset of $\mathsf{var}(B)$ so that the following holds. Let $P$ be a directed path of $B$ and suppose that $P$ has two decision nodes $u_{1}$ and $u_{2}$ labelled by variables $x_{1}$ and $x_{2}$ and assume that $u_{1}$ occurs before $u_{2}$ in $P$. Then $x_{1}<_{\pi}x_{2}$ (that is $x_{1}$ is smaller than $x_{2}$ according to $\pi$). Then $B$ is called an $\wedge_{d}$-obdd (respecting $\pi$ if not clear from the context). Note that the $\wedge_{d}$-fbdd $B$ considered in Example 1 is in fact an $\wedge_{d}$-obdd respecting an arbitrary order over $\{x,y_{1},y_{2},y_{3}\}$ provided that $x$ is ordered first. Finally, an $\wedge_{d}$-obdd without $\wedge_{d}$ nodes is just an obdd.
We remark that allowing $\pi$ to be defined over a superset of $\mathsf{var}(B)$ will be convenient for our reasoning when we consider subgraphs of a $\wedge_{d}$-obdd.
@@
Throughout the paper, when we refer to the size of a smallest model representing a function $f$, we will use the notation $modelname(f)$. For instance $\wedge_{d}$-obdd(f)(f), fbdd(f)(f), obdd(f)(f) respectively refer to the smallest sizes of $\wedge_{d}$-obdd, fbdd, and obdd representing $f$.
@@ 2.3 Graphs, width parameters, and correspondence with CNF
We use the standard terminology related to graph as specified in [9]. In particular for a graph $G$, $U\subseteq V(G)$ and $E^{\prime}\subseteq E(G)$, $G[U]$ and $G[E^{\prime}]$ respectively denote the subgraphs of $G$ induced by $U$ and $E^{\prime}$ Also, let us recall that a matching $M$ of a graph $G$ is a subset of its edges such that for any distinct $e_{1},e_{2}\in M$, $e_{1}\cap e_{2}=\emptyset$. The treewidth of $G$, denoted by $tw(G)$ is arguably the best known graph parameter measuring closeness to a tree. To define $tw(G)$, we need the notion of a tree decomposition of $G$ which is a pair $(T,{\bf B})$ where $T$ is a tree and ${\bf B}:V(T)\rightarrow 2^{V(G)}$. The set ${\bf B}(t)$ for $t\in V(T)$ is called the bag of $T$. The bags obey the rules of containement meaning that for each $e\in V(G)$ there is $t\in V(T)$ such that $e\subseteq{\bf B}(t)$ and connectivity. For the latter, let $T[v]$ be the subgraph of $T$ induced by $t$ such that $v\in{\bf B}(t)$. Then $T[v]$ is connected. For $t\in V(T)$, the width of ${\bf B}(t)$ is $|{\bf B}(t)|-1$ The width of $(T,{\bf B})$ is the largest width of a bag. Finally, $tw(G)$ is the smallest width of a tree decomposition. The pathwidth of $G$, denoted by $pw(G)$, is the smallest width of a tree decomposition $(T,{\bf B})$ of $G$ such that $T$ is a path.
@@
For a cnf $\varphi$, we use two graphs, both regard $\varphi$ as a hypergraph over $\mathsf{var}(\varphi)$ ignoring the literals with clauses corresponding to hyperedges. The primal graph of $\varphi$ has $\mathsf{var}(\varphi)$ as the set of vertices and the two vertices are adjacent if they occur in the same clause. The incidence graph of $\varphi$ has $\mathsf{var}(\varphi)$ and the clauses as its vertices. A variable $x$ is adjacent to a clause $C$ if and only if $x$ occurs in $C$. We denote by $ptw(\varphi)$ and $itw(\varphi)$ the trewidth of the primal and incidence graphs of $\varphi$ respectively.
@@
3 $\wedge_{d}$-OBDD: inefficiency in terms of incidence treewidth and separations
3.1 An approach to proving $\wedge_{d}$-OBDD lower bounds
We start from overviewing an approach for proving lower bounds for obdds and then generalize it to $\wedge_{d}$-obdds. This is the standard ‘lower bound by the number of subfunctions’ approach as described in [25] but rephrased for the purpose of the further generalization.
Let $B$ be an obdd respecting an order $\pi$. Let $\bf g$} be an assignment over a prefix of $\pi$. Then we can identify the maximal path $P=P({\bf g})$ starting from the source and such that ${\bf a}(P)\subseteq{\bf g}$. Indeed, existence of such a path and its uniqueness is easy to verify algorithmically. Let the current vertex $u$ of $B$ be the source and the current variable $x$ be the label of $u$. Furthermore, let the current path $P$ be consisting of $u$ only. If $x\notin\mathsf{var}({\bf g})$ then stop. Otherwise, choose the out-neighbour $v$ of $u$ such that $(u,v)$ is labelled with ${\bf g}(x)$. Let the current vertex be $v$, the current variable be the label of $v$ and the current path be $P$ plus $(u,v)$. Repeat the process until the stopping condition reached. We let $u({\bf g})$ be the final vertex of $P({\bf g})$.
Next, we define $\mathcal{F}$, a set of fooling assignments, for each $\bf g$$\in\mathcal{F}$, $\mathsf{var}({\bf g})$ is a prefix of $\pi$. The idea is to prove that for every two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$. This implies that $|B|\geq|\mathcal{F}|$. In particular, if $|\mathcal{F}|$ is large in terms of the number of variables, then the respective large lower bound follows.
For $\wedge_{d}$-obdds, an assignment to a prefix of the respected order is associated with a rooted tree rather than with a path. In the rest of the subsection, we upgrade the above methodology for this case.
@@ 3.1 An approach to proving $\wedge_{d}$-OBDD lower bounds
Figure 2 demonstrates a $\wedge_{d}$-obdd $B$ for $(x_{1}\vee x_{2})\wedge(x_{2}\vee x_{3})\wedge(x_{4}\vee x_{5})\wedge(x_{5}\vee x_{6})$ (on the left) respecting the order $(x_{2},x_{1},x_{3},x_{5},x_{4},x_{6})$ and $B[{\bf g}]$ (on the right) where ${\bf g}=\{(x_{2},1)\}$.
@@
A decision node of $B[{\bf g}]$ that has only one outgoing neighbour is called an incomplete decision node. A maximal path with all intermediate decision nodes being incomplete is called an incomplete path. We note that a $\wedge_{d}$ node cannot be a final node of an incomplete path due to its maximality, hence an incomplete path may end only with a sink of $B$ or with a decision node of $B$. We denote by $L({\bf g})$ the set of all decision nodes that are final nodes of the incomplete paths of $B[{\bf g}]$.
Back to Example 3, $L({\bf g})$ is a singleton consisting of the only node labelled by $x_{2}$. We note that $L({\bf g})$ are exactly minimal complete decision nodes of $B[{\bf g}]$ in the sense that each decision node that is a proper descendant of a node of $L({\bf g})$ must be complete.
@@
Let $\pi$ be the order respected by $B$ and let $\bf g$} be an assignment over a prefix of $\pi$. Then the following statements hold regarding $B[{\bf g}]$.
@@
    Let $T({\bf g})$ be the union of all incomplete paths ending with nodes of $L({\bf g})$. Then $T({\bf g})$ is a rooted tree. Moreover, all the intermediate nodes of $T({\bf g})$ with out-degree $2$ are conjunction nodes. In particular, this implies that if $B$ is an obdd then $T({\bf g})$ is just a path.
@@
Let $x$ and $y$ be the variables labelling $u$ and $v$ respectively. By assumption $x<_{\pi}y$ and $x\notin\mathsf{var}({\bf g})$, the latter constitutes a prefix of $\pi$. It follows that $y\notin\mathsf{var}({\bf g})$ and hence $v$ is complete simply by construction.
@@
For the third statement, it is sufficient to prove that any two distinct incomplete paths $P_{1}$ and $P_{2}$ do not have nodes in common besides their comon prefix and that the last node of the prefix is $\wedge_{d}$ one. So, let $P$ be the longest common prefix of $P_{1}$ and $P_{2}$ and let $u$ be the last vertex of $P$. Arguing as in the previous paragraph, we conclude that $u$ is a $\wedge_{d}$ node. Let $u_{1}$ and $u_{2}$ be the children of $u$. Then, by the maximality of $P$, we can assume w.l.o.g. that $u_{1}\in V(P_{1})$ and $u_{2}\in V(P_{2})$.
@@
Lemma 2 is illustrated in Figure 3. Suppose that the picture illustrates $B[{\bf g}]$ for some $\wedge_{d}$-obdd $B$ and an assignment $\bf g$} over a prefix of the order respected by $B$. Then the rooted tree within the outer dashed rectangle is $T({\bf g})$, the set of nodes in the inner dashed is $L({\bf g})$. We note that the nodes of $T({\bf g})$ with two children are the $\wedge_{d}$ ones. Each $u\in L({\bf g})$ is contained in an ellipse and the ellipse denotes $B_{u}$, this is demonstrated for the leftmost ellipse. All the $\wedge_{d}$-obdd are assumed to have the same $True$ and $False$ sink, we do not demonstrate this on the picture for the sake of clarity.
$uB_{u}$$L({\bf g})$$T({\bf g})$$\wedge_{d}$$\wedge_{d}$
Figure 3: An intuitive illustration of $T({\bf g})$ and $L({\bf g})$
The notion of $L({\bf g})$ allows a neat characterization of $\mathcal{S}(B)|_{\bf g}$ as specified in the next lemma.